# Review, Refine, Repeat: Understanding Iterative Decoding of AI Agents with Dynamic Evaluation and Selection

### v1

@@ Abstract
While AI agents have shown remarkable performance at various tasks, they still struggle with complex multi-modal applications, structured generation and strategic planning. Improvements via standard fine-tuning is often impractical, as solving agentic tasks usually relies on black box API access without control over model parameters. Inference-time methods such as Best-of-N (BON) sampling offer a simple yet effective alternative to improve performance. However, BON lacks iterative feedback integration mechanism. Hence, we propose Iterative Agent Decoding (IAD) which combines iterative refinement with dynamic candidate evaluation and selection guided by a verifier. IAD differs in how feedback is designed and integrated, specifically optimized to extract maximal signal from reward scores. We conduct a detailed comparison of baselines across key metrics on Sketch2Code, Text2SQL, and Webshop where IAD consistently outperforms baselines, achieving 3–6% absolute gains on Sketch2Code and Text2SQL (with and without LLM judges) and 8–10% gains on Webshop across multiple metrics. To better understand the source of IAD’s gains, we perform controlled experiments to disentangle the effect of adaptive feedback from stochastic sampling, and find that IAD’s improvements are primarily driven by verifier-guided refinement, not merely sampling diversity. We also show that both IAD and BON exhibit inference-time scaling with increased compute when guided by an optimal verifier. Our analysis highlights the critical role of verifier quality in effective inference-time optimization and examines the impact of noisy and sparse rewards on scaling behavior. Together, these findings offer key insights into the trade-offs and principles of effective inference-time optimization.
@@ 1 Introduction
AI agents have demonstrated remarkable capabilities across a wide range of tasks, including natural language understanding, code generation, and mathematical reasoning (Fu et al., 2024; Liu et al., 2024; Jiang et al., 2024; Patel et al., 2024; Huang et al., 2024; Liu et al., 2023). However, state-of-the-art AI agents still face challenges when handling tasks that require understanding multimodal inputs and reasoning through structured output generation (Costarelli et al., 2024; Kinniment et al., 2024; Liu et al., 2023; Verma et al., 2024; Valmeekam et al., 2024). For example, consider agentic tasks such as Sketch2Code (Li et al., 2024b), Text2SQL (Li et al., 2024a), and the web-navigation task Webshop (Yao et al., 2023). Sketch2Code requires an AI agent to generate structured HTML from rough sketches, often resulting in incorrect generations. Text2SQL involves translating natural language queries into executable SQL statements, demanding precise reasoning while avoiding semantic and logical errors. Meanwhile, Webshop is a sequential decision-making task that evaluates an agent’s ability to navigate an e-commerce platform effectively. Across these tasks, AI agents consistently underperform compared to human experts, underscoring the limitations of current agentic models in handling complex, structured outputs and sequential reasoning (Yao et al., 2023; Liu et al., 2024; Li et al., 2024b) where agents achieve merely 20-30% performance in task specific metrics.
Figure 1: Qualitative (Left) and Quantitative (Right) illustration of the performance benefit of IAD (Ours) over Single-turn response generations using Gemini-1.5 (Base) Text2SQL task (Li et al., 2024a). IAD improves performance by correctly handling query logic and joins, improving the accuracy over baseline and BON (Best-of-N).
A key challenge in improving agentic AI systems is their ability to handle structured reasoning and multimodal tasks effectively. One promising approach to addressing this challenge is fine-tuning a new model through alignment (Ouyang et al., 2022a). A key alignment technique, reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022a, b; Bai et al., 2022) has significantly enhanced the capabilities of generative models, it is not directly applicable to AI agents due to their inherently black box nature. Most agentic systems and frameworks (Liu et al., 2024; Fu et al., 2024) operate in environments where internal model updates are infeasible, making fine-tuning-based alignment techniques inapplicable. An alternative approach is inference-time alignment, such as controlled decoding (Khanov et al., 2024; Deng and Raffel, 2023; Mudgal et al., 2024; Han et al., 2024; Chakraborty et al., 2024; Yang and Klein, 2021), which can align generative models without fine-tuning. However, these methods still rely on access to token-level logits from the reference model to steer responses, an assumption that does not hold for black box agentic systems. These limitations highlight the need for an inference-time alignment technique that is (i) compatible with various underlying models (without requiring model-specific adaptations), (ii) easy to implement (requires no modification of the underlying models), and (iii) commercial API-friendly (Applicable to closed-source models that are only accessible via API. This leads us to an important question:
How can we improve the inference alignment of AI Agents with black box access for complex and structured tasks?
A simple solution to the above question is Best-of-N (BON) sampling and its variants (Nakano et al., 2021; Stiennon et al., 2020; Beirami et al., 2024), as it is black box compatible and logit-free, making it a natural choice for inference-time alignment. BON selects the best response from multiple generated candidates using a verifier, improving response quality in structured tasks such as HTML generation and SQL query synthesis. BON is even almost optimal in terms of quality (Gui et al., 2025; Yang et al., 2024b; Beirami et al., 2024; Amini et al., 2024). However, BON may be ineffcient requiring a large N to achieve reasonable performance. To address these limitations, we propose Iterative Agent Decoding (IAD)—an iterative approach that refines responses through verifier-guided feedback. IAD differs fundamentally in the design and integration of its feedback mechanism, specifically optimized to extract and utilize rich signals from the reward score thus demonstrating the effectiveness of adaptive feedback in complex tasks. We then conduct a detailed comparison of baselines including parallel sampling (BON) and iterative decoding (IAD) across several key metrics relevant to inference-time scaling, yielding key insights into the trade-offs and principles of effective inference-time optimization. We summarize our contributions and findings as follows.
Figure 2: Demonstrates the high-level schematic of our proposed approach - Iterative Decoding of AI Agents
1. An iterative decoding framework (IAD): We propose Iterative Agent Decoding (IAD), an efficient iterative framework where responses are refined through verifier-guided feedback, enabling improved performance in black box structured generation tasks. IAD is designed to extract maximal information from available feedback, including sparse or noisy scores.
2. Empirical evaluation across agentic tasks: We evaluate IAD on three complex structured output tasks, demonstrating consistent improvements over strong baselines, while maintaining efficiency. IAD minimizes reliance on crafted exemplars and leverages feedback-driven reconditioning, making it task-agnostic and model-agnostic.
3. Disentangling feedback gains in IAD from sampling: We conduct controlled experiment to isolate the effect of adaptive feedback in IAD from sampling diversity under low-temperature sampling, where IAD continues to improve, while BON performance stagnates, confirming that IAD’s gains are driven by adaptive, feedback-guided refinement, not sampling diversity alone.
4. Inference scaling behavior: We highlight the inference-time scaling of performance improvements where we show that both IAD and BON improve with increased test-time compute when guided by an optimal verifier. However, IAD benefits more significantly from high-quality feedback, leading to greater performance gains. Additionally, as model capability increases, BON performance improves, narrowing the gap with IAD.
5. Importance of verifier quality for inference scaling: We show that high-quality verifier feedback is critical for inference scaling. Self-evaluation or sub-optimal verifiers often fail in complex agentic scenarios, whereas optimal verifier-guided alignment enables consistent refinement.
6. Robustness under sparse and noisy feedback: To further illustrate the impact of verifier quality, we conduct controlled experiments by sparsifying and adding noise to the verifier signal. Results show that IAD remains robust under mild sparsity and noise, but performance for both IAD and BON degrades under extreme conditions, due to the lack of reliable feedback.
Remarks: Several multi-turn/multi-attempt iterative approaches focus on train-time methods (Qu et al., 2024; Zhou et al., 2024; Shani et al., 2024), making them unsuitable for black box inference. Inference-time methods (Madaan et al., 2023; Chao et al., 2024; Mehrabi et al., 2024; Muennighoff et al., 2025) have explored iterative refinement in specific contexts like jailbreak, math, and QA, but primarily for simpler tasks rather than complex agentic settings requiring structured generations like HTML and SQL. Specifically, Chao et al. (2024); Mehrabi et al. (2024) iteratively refine prompts for adversarial attacks. Meanwhile, Madaan et al. (2023) and its variants emphasize on LLM self-evaluation rather than focusing on the design choices for iterative refinement based approach for structured output generation.
2 AI Agent Alignment
In this section, we formalize the problem of inference-time alignment for AI agents operating in black box settings, particularly in complex structured-generation tasks such as HTML and SQL synthesis in multi-modal environments like Sketch2Code, Text2SQL, and Webshop. Given a task distribution $\mathcal{T}\in\Delta$, we define the optimal but inaccessible agent policy $\pi^{*}(\cdot|x)$, which generates the ideal structured response $y^{*}\sim\pi^{*}(\cdot|x)$ for a given input prompt $x$. However, in the black box setting, we only have access to a reference policy $\pi_{0}(\cdot|x)$, which may lack some capabilities needed to complete the task.
Figure 3: Sketch2code: This figure provides a comparison of IAD (ours) against Best-of-N sampling (SoTA) and single-turn generation with Gemini-1.5-Pro with temperature =0.6 w.r.t metrics - (a) Layout Similarity (b) TextIoU (c) ImageIoU across varying the number of generations (N). Figure demonstrates IAD outperforms BON consistently across N, but as N increases the gap reduces. Note that (b) and (c) are not directly optimized for this plot and test the generalization performance. This demonstrates the consistency in performance across models and metrics

### v4

# On the Role of Feedback in Test-Time Scaling of Agentic AI Workflows

@@ Abstract
Agentic AI workflows (systems that autonomously plan and act) are becoming widespread, yet their task-success rate on complex tasks remains low. A promising solution is inference-time alignment, which uses extra compute at test time to improve performance. Inference-time alignment relies on three components: sampling, evaluation, and feedback. While most prior work studies sampling and automatic evaluation, feedback remains under-explored. To study the role of feedback, we introduce Iterative Agent Decoding (IAD), a procedure that repeatedly inserts feedback extracted from different forms of critiques (reward models or AI-generated textual feedback) between decoding steps. Through IAD, we analyze feedback along four dimensions: (1) its role in the accuracy–compute trade-offs with limited inference budget, (2) quantifying the gains over diversity-only baselines such as best-of-N sampling, (3) effectiveness of composing feedback from reward models versus textual critique, and (4) robustness to noisy or low-quality feedback. Across Sketch2Code, Text2SQL, Intercode, and WebShop, we show that IAD with proper integration of high fidelity feedback leads to consistent gains up to 10% absolute performance improvement over various baselines such as best-of-N. Our findings underscore feedback as a crucial knob for inference-time alignment of agentic AI workflows with limited inference budget.
@@ 1 Introduction
Figure 1: General pipeline for inference-alignment approaches for AI agents.
AI agents demonstrate remarkable potential across a wide range of applications, yet they still face challenges in complex tasks requiring strategic planning and multimodal reasoning. For example, on challenging benchmarks such as Sketch2Code (Li et al., 2024b), Text2SQL (Li et al., 2024a), Intercode (Yang et al., 2023a), and Webshop (Yao et al., 2023), even state-of-the-art agents can only achieve 20–30% accuracy. While post-training via SFT and RL (Ouyang et al., 2022a, b; Bai et al., 2022) has been an effective tool to enhance the capabilities of generative models; however, it is not directly applicable to AI agents. This is because AI agents frequently operate as black-box systems with no access to the internals (Liu et al., 2024; Fu et al., 2024). Hence, we focus on inference-time approaches which are (i) API-friendly (applicable to closed-source models accessible via commercial APIs) and (ii) compatible with various underlying models.
At inference time, alignment strategies broadly iterate over three components: sampling, evaluation, and feedback as illustrated in Figure 1. A typical inference-time iteration involves sampling one or more outcomes from the generative model, evaluating them using a judge/reward, and then generating feedback to improve subsequent outcomes. The role of sampling and evaluation is well-studied in literature with efficient and scalable algorithms based on best-of-N (BoN) approaches (Nakano et al., 2021; Beirami et al., 2024). BoN-based approaches operate by sampling multiple responses and selecting the best according to a verifier, effectively leveraging the sampling and evaluation steps. While BoN and variants (Nakano et al., 2021; Beirami et al., 2024; Verdun et al., 2025) are highly parallelizable, allowing responses to be sampled and evaluated concurrently to reduce wall-clock latency, it may not be compute-optimal. That is, to achieve a desirable outcome, BoN may require generating a large number of samples and evaluating each independently. This parallelism trades off compute efficiency for lower latency. Therefore, when compute is constrained, simply generating more samples is not a viable path to better performance, highlighting the need for a more adaptive approach. More importantly, a key limitation of BoN approaches is their inability to incorporate feedback mechanisms for iterative refinement.
To address the above-mentioned limitations, recently, several sequential approaches have emerged (Madaan et al., 2023; Yang et al., 2023b; Chen et al., 2023; Robeyns et al., 2025; Jiang et al., 2024) that aim to explicitly integrate feedback for iterative refinement. This trend towards self-correction is evident in techniques such as self-refine (Madaan et al., 2023), where models iteratively improve their outputs based on self-critique, and try-again strategies (Yang et al., 2023b) that prompt the LLM to generate alternative responses when initial attempts fail. Other approaches include (Chen et al., 2023), where LLMs identify and correct errors in their own generated code. Additionally, recent advancements in agentic systems have showcased their ability to self-improve (Robeyns et al., 2025; Jiang et al., 2024).
Figure 2: Overview of Iterative Agent Decoding (IAD), highlighting the sampling, evaluation, and feedback components of inference-time alignment. The figure illustrates how feedback is constructed from verification signals and used to iteratively refine model outputs.
While existing methods creatively use feedback, they do not systematically study it. They often lack a focused analysis of how feedback should be designed, integrated, or optimized. While prior approaches (Chen et al., 2023; Yang et al., 2023b; Madaan et al., 2023) demonstrate the utility of adaptive learning with feedback, a clear understanding of the mechanisms by which LLMs learn from this feedback, especially with scalar feedback or score, is still developing. As a result, the role of feedback in inference-time alignment, especially in black-box agentic environments, remains under-examined. We aim to address this gap by focusing on feedback and systematically understanding its impact on test-time scaling of agentic workflows. Specifically, we focus on inference-time alignment strategies designed to maximize performance under a fixed compute budget, where compute is defined concretely as the total number of API calls or forward passes through the model. In such compute-constrained settings, incorporating feedback becomes critical for effective test-time scaling. Thus, while BoN strikes a balance between latency and compute, approaches that leverage feedback are better suited when compute is the principal focus.
Our contributions (Understanding the role of feedback). To understand the role and impact of feedback on inference-time alignment, we require a unified framework that flexibly integrates diverse forms of feedback, guided by a verifier, without relying heavily on prompt engineering, which can be costly and fragile. To this end, we introduce Iterative Agent Decoding (IAD), a sequential framework designed to incorporate various forms of feedback to understand the role of feedback in inference-time alignment of agents. Through IAD we uncover several intriguing results, detailed below.
1. Accuracy vs compute: budget-constrained scaling with feedback: We demonstrate that feedback plays a critical role in budget-constrained settings and helps achieve up to a 10% accuracy gain over strong feedback-free baselines with the same budget. However, this margin narrows with increased budget, which allows for the generation of more candidates (e.g., larger $N$ in BoN) or using more capable models like Gemini-1.5 to Gemini-2.0 and Gemini-2.5 (cf. Figure 3).
2. Impact of adaptive feedback beyond sampling gain: We perform controlled experiments to capture the gains from feedback beyond the diversity gains from sampling, and demonstrate that the gains in IAD arise from adaptive, feedback-guided refinement, yielding performance improvements of 4–8% in both Sketch2Code and Text2SQL. This also highlights that feedback plays a crucial role in agentic tasks and scenarios where sampling diversity is low.
3. Design and role of feedback form: We investigate the impact of different forms of feedback, primarily scalar and textual on inference-time alignment. Textual feedback is relatively straightforward to integrate, as it can be directly incorporated into the model prompt. However, designing effective feedback from scalar rewards and preferences remains underexplored. In IAD, we focus on methods to extract useful signals from scalar rewards and demonstrate that the performance gains crucially depend on the design of proper feedback. Our experiments show that transforming scalar feedback into directional prompts yields gains of up to 6–7% over baselines in Sketch2Code and Text2SQL.
4. Sensitivity to feedback quality: We study the sensitivity and robustness of inference alignment to feedback quality via controlled sparsity and noise in feedback signals. We see that IAD remains effective under moderate degradation, but its performance declines with increasing noise. Under high sparsity or heavy noise, the performance gains over sampling-based methods drop by approximately 4–5%, highlighting the importance of feedback fidelity for reliable test-time scaling.
2 Inference-time Agent Alignment